Preparation visit before Run6
In May 2013, 3 people come to the 30m to install the new 2mm array and replace the 1mm cold amplifier. Then, a cooling down will occur and the NIKA prototype stays cold until the Run6 proper starts.
Tuesday 14-May-2013: 10h (during maintenance) Open cryostat, replace array at 2mm and the 1mm cold amplifier. Install the new DC cables for the amplifier. Install a new 4K magnetic screen. Install electronics rack in the cabin and connect. Mount the anti-vibration legs on the pumping stage. In the evening, start cooling down.
15-17 May 2013: cooling down
18 May 2013: check resonances in a dark configuration and end the preparation as soon as it is ready. The cryostat stays cold until the observational run 6 in June.
Alessandro Monfardini, Martino Calvo and Nello D'Addabbo come by car from Grenoble.
PreparationRun Sun 12/05
Left Grenoble by car in the morning, stopped at Calp
PreparationRun Mon 13/05
Arrived at the telescope around 3:30pm. Unloaded the car, installed archeops8, archeops9 and THz in the control room. Updated with SVN. Experiment file for next NIKA run: nika_iram2.txt
A few mbars of something in the cryostat (J1=50mbar).
PreparationRun Tue 14/05
GISMO electronics is still there. Thus first we'll have to wait that it's craned down before starting working on the cryostat. We re-pump the cryostat and the big pumping pipe before the intervention. Jpi is growing up a bit after pumping; 1e-3mbar per minute roughly. Not very dangerous, could be outgassing, but I have seen better. Could be a small leak at the O-rings. In any case we have to open later.
All planned operations completed. Furthermore, mu-metal screening on 4K and 50K stages:
- on 4K screen: (inner) METGLAS 2714A 2.0 inch x 0.6 mil std spec (bat), mu about 10000 @ 4K (see figure below); (outer) VITROVAC (same put on Sionludi @ Institut Néel)
- on 50K screen: only METGLAS
figure: METGLAS mu at 4.2K, real and imaginary parts.
Installed NICA_8f1 18nm IEF (litho at PTA) tested in run HE34 in Grenoble at 2mm. At 1.25mm left the NICA_8e 20nm IRAM we had for runs 4 and 5.
Cooling down started at 17:30. Everything seems correct. Only (known already) problem is a small leak (1.5e-7) while circulating for the preref. The leak is proportional to the flux.
PreparationRun Wed 15/05
Fixed a problem preventing the NIKEL boards to kick on automatically (still to be investigated in Grenoble with Olivier). Fixed an annoying problem preventing to control properly the cryostat. We have installed a brand new blue ethernet cable going directly from the main NIKA switch in the cabin down to the "automate" of the cryostat. Now the connection is OK, no more risks of losing control so far.
In the afternoon, taking advantage of bad weather preventing observations, we have installed the whole NIKA optics (curved M6) and laser-aligned the system in order to be ready for the next run. Replaced, with Santiago and Gregorio, one broken anti-vibration leg of the table holding the cryostat. We also made new semi-rigid cables for better connecting the electronics to the cryostat. Gregorio says we have used the VERY LAST one available at Pico Veleta. Since they break very often (every time more or less) not to be forgotten buying few more before the next run.
21h30. Snowing. T_4K=6K; T_still=8K; T_BM=9K. Reduced strongly the preref flux. Pumping through the 3He pump, injection pressure J4=1200mbar, on the pump input 15mbar (should be safe). Leak signal is now 3e-8. Refilled LN2 trap.
figures: Cryostat status at 21h30 today. J1 zero-point is at -15mbar actually, means the 3He pump in is at 15mbar.
PreparationRun Thu 16/05
Problem during the night: the "rapide" circulation was not correctly launched and when the first drops of He started to condense at the 4K coldest part the flux dropped drammatically. Thus we've lost roughly 6 hours of cooling down. In the morning, put again the "rapide" circulation" and started condensation at 10am. Not working immediately (still and BM heating up); had to open ev18 to have it started. At 14h everything seems OK. Still and BM are cooling toward 4K and the injection pressure J4 drops.
Procedure adopted this time for cooling down the cryostat (DO NOT USE IN GENERAL, EVERY TIME IS DIFFERENT):
1) "Rapide" (ev5=1; ev2=1; ev1=ev17=ev19=0) with J2 initially at 750mbar (J4 around 3500mbar), going down slowly. Continue this way until T_still = T_4K (usually satisfied at about 20K; takes roughly 16h).
2) J2=500mbar (ev5=1) until further flattening; then J2=300mbar (ev5=1) until further flattening.
3) When (T_still < 10K and T_BM < 10K), close ev5, ev5=0. Steps 2 and 3 together take about 4-6 hours. From now on ev5 stays closed.
4) When flattening again, the only other choice is passing through the pump (ev2=0, ev1=ev17=1, compressor ON, J1=15-20mbar); now it should be 24h roughly from the beginning. Pay attention to the J1 zero-point, could be not zero.
5) Leave this way (ev2=ev5=0; ev1=ev17=ev19=ev3=ev6=1; ev7=ev8=1; preref ON regulating J2 at 300mbar; compressor ON) for the night or in any case for about 8 hours until (T_still < 7K and R_BM > 1240-Ohm). If they're still cooling down reasonably when you arrive there, you might leave further until flattening. Very important to leave the "preref" switched ON. Otherwise, when the 4K reaches the LHe temperature a large part of the gas circulating condenses, leaving with a largely insufficient gas flux to cool down. So you basically lose the night entirely.
6) If (T_still < 7K and R_BM > 1240-Ohm) you could try to condense the mixture. To do that, simply close ev7, thus ev7=0, and switch ON the condensation procedure to regulate J4 at 4000mbar. Open ev18 to help the still and BM cooling. In case, increase slightly the pressure to J4=4500mbar (never more than that) but then decrease the target regulation value to 4000mbar.
7) At some point at the beginning of the condensation (e.g. after a few cycles of ev11 open/close) you might want to close ev18 to open it again when J5<50mbar. Or not, it depends on how it goes. Condensation should take no more than 12 hours (TBC).
8) When (J5 < 10mbar and J4 < 2000mbar) open ev18 if it's closed. ev18=1. Switch ON the turbo and stop the condensation procedure. Now just a matter of waiting it to reach base T.
At 19h the condensation is going well. J5=160mbar, going down at a rate of 40mbar/hour. So tomorrow morning it should be cold.
Availability of planets for Sunday 6-8 UTC slot this horizon plot
22h30. BOTH ARRAYS ARE CONNECTED !! With the amplifiers ON the overall gain of the cryostat is about +10dB (-20dB attenuation - 5dB cables +35dB amplifier gain) as expected. The condensations has finished but the BM is still warm at 1.2K. We have connected the semi-rigid cables TO/FROM the electronics so tomorrow we don't have to go to the cabin.
PreparationRun Fri 17/05
During the night, the Turbo was ON and the cryostat cooled slowly down to 250mK, reached at about 8h in the morning.
The resonances on the 2mm array are very nice since the beginning. 129 good resonances, noise around 1Hz/sqrt(Hz) (but please consider the window on the cryostat is closed). No glitches. Exactly like in lab. The amplifier on the 1mm channel seems not saturating even with -50dBm per tone at input and 300 tones. At the output we have -36dBm per tone, meaning that the amplifier has the correct gain.
The resonances of the 1mm, even seen on the VNA, are VERY small (<<1dB) and large (MHz!). This could be explained if the array was still warm, say at about 400mK. Or if the lens in front of it is not thermalized yet. That's a problem we had recently in the lab on the He cryostat. In any case it could be possible that a "hot spot" near the 1mm array heats this last but not much the 2mm. Nothing is lost since we have checked the last cooldown plots and seen that the cryostat had done the same for about 18h before cooling properly at some point at <100mK. So we decide to wait until tonight and see if it does the same this time. The last time we were in remote from Granada following the cooldown during the GISMO run, so we don't know if the resonances at that time were there or not. Optimistically we could hope that tonight the cryostat behaves like the last two times.
ACTION for the NEXT TECHNICAL INTERVENTION (e.g. WHEN WE'LL CHANGE THE 1.25mm ARRAY ?): CHECK CAREFULLY THE MOUNTING OF THE HDPE LENSES AT 100mK. THEY REALLY NEED TO BE SQUEEZED VERY TIGHTLY. LESSON LEARNED FROM GRENOBLE CRYOSTAT AND EDELWEISS.
Since the cryostat seems in trouble going down (keeps oscillating between 240 and 270mK) and the resonances on the 1.25mm array are getting better very slowly we decide to open up and check immediately. At 19h30 we start heating up: stopped the PT, fast 4He circulation, a lot of exchange gas (LN2). At 22h we're in the cabin to start opening up. One screen every 20 minutes roughly. We decide not to move the cryostat and do the intervention in the nominal position. Just we switch off the anti-vibrations pistons. IT WAS FINALLY POSSIBLE, even if acrobatic, making things a bit easier than thought. We reached the 100mK stage at 00h30 and seen with surprise that the lens at 1.25mm was well thermalized. On the other hand we have found (and removed) 10 loosly screwed big black paint plaques at 100mK (mounted well behind the 1.25mm array, so not THAT useful), two M4 stainless steel bolts floating at the bottom of the still box, just in front of the 100mK and one screw holding the dichroic without any bolt, just standing there with no thermalization at all.
PreparationRun Sat 18/05
We finish closing down at about 2h00 in the morning, everything looks fine. We pump down, no leaks at 2h30. BUT WE RALIZE THAT THE DILUTION INJECTION IS BLOCKED. We try many things until 3h30 in the morning (Alain joined us with Skype at 3h00). There's nothing we can do without opening the cryostat. There's still hope that the injection is blocked by something left cold after the fast heat-up. So we decide to go sleep for a few hours leaving 4bars at the injection. Hoping to wake up and found it circulating. 8h: after just 4 hours resting (also, quite agitated) we wake up and realize it's still blocked. So we have to open the cryostat, unmount the optics, the arrays, the 100mK plate (this last step is something we do very rarely even in Grenoble). We move the pumping stage with the cryostat on top in the corridor away from the beam. We open up and realize that it's the first injection impedance that's blocked. We decide to go for lunch in any case, we need forces. 14h: after two-three tries, we cut the impedance in half and we found one half that's OK !! The problem has been isolated. Now it seems we just have to remount the impedance on the capillary going to the still and that's it !! Unfortunately at this point we realize that the capillary going to the still is blocked too and we don't understand why. Between 14h and 16h we try unblocking it with no success. Then we realize the problem: one manual valve had been closed on the pumping stage to allow safer moving of the cryostat out of the beam. And then forgotten. That means when we unsoldered the injection impedance we had unexpectedly (on our screen we saw differently, couldn't see the manual valve) low pressure on the still side. So the soldering has been sucked into the still capillary. Shit. How deep has it penetrated into the still ?? Nodoby knows. We continue until 20h, heating the capillary and entering very carefully with small Cu wires to clean it. Despite the efforts, still blocked at 19h30. We got rid of a lot of Sn in the capillary but it became the more and more difficult the more we entered into the still. We couldn't heat enough and put enough pressure to unblock it. At 20h, after a Skype discussion with Alain, we decide we have to decide something in any case. Our decision was to bring the cryostat back to Grenoble AS SOON AS POSSIBLE to repair it. We prepare the transportation between 21h and 22h, then we crane it down to the garage between 22h and 23h with the help of Salvador and Victor. Between 23h and 24h we clean a bit the cabin that was really a huge mess. We are tired.
PreparationRun Sun 19/05
Sunshine, the telescope is finally moving. It was a good decision to crane down the cryostat yesterday. We sleep a bit more, after the two very difficult days spent in the cabin below 0C. In the morning we decided to finish up adding magnetic screening bu taping also the 150K screen. So now we have the following:
- 300K: mu-metal 2mm
- 150K: Metglas 2714A 0.6mil (protected by Al-Mylar)
- 50K: Metglas 2714A 0.6mil (protected by Al-Mylar)
- 4K: Metglas 2714A 0.6mil + one Vitrovac sheet
- 1K: Pb very thick superconducting screen
In the afternoon we prepare everything for tomorrow. Remounted the screens in the garage and pack the stuff. The car is at the National Park gate and there's still a lot of snow, so it won't be smooth. Carsten, Santiago and everybody else here are working to make it possible, we'll see tomorrow. List of things we leave at Pico Veleta:
NIKEL01 and NIKEL02 mounted in the rack (receivers cabin). Leaving the rack switched on and connected for remote electronics tests. Shorted IN/OUT, with 3dB attenuation on NIKEL02 and nothing on NIKEL01.
N.4 Semi-rigid cables we had prepared with the good shape (receivers cabin)
Anti-vibration legs for the pumping stage (receivers cabin)
Dilution pumping stage (receivers cabin)
M5 and M6 with their mountings (receivers cabin)
DN40 pipe connecting the cryostat and the pumping stage (receivers cabin)
"THz" (192.168.1.10), "archeops8" (192.168.1.8), "archeops9" (192.168.1.9) (control room). Switched on and connected to the NIKA network.
"Windows cryostat" (192.168.1.31), "Linux analysis" (192.168.1.32), "Mini-Mac acquisition" (192.168.1.33), Router (192.168.1.3) (computers room). Switched on and connected to the NIKA network.
Ethernet cables, power cables, mouses, small computer equipment (computers room).